# Smooth orbit equivalence rigidity for dissipative geodesic flows

### v1

@@ Abstract.
Let $M$ be a smooth closed orientable surface. A Gaussian thermostat on $M$ can be seen as the geodesic flow of a certain metric connection with torsion. These flows may not preserve any smooth volume form. We prove that if two Gaussian thermostats on $M$ with negative thermostat curvature are related by a smooth orbit equivalence isotopic to the identity, then the two background metrics are conformally equivalent via a smooth diffeomorphism of $M$ isotopic to the identity. We also give a relationship between the thermostat forms themselves. Finally, we prove the same result for Anosov magnetic flows.
@@ 1. Introduction
Let $(M,g)$ be a smooth closed oriented Riemannian surface, and let $\lambda\in\mathcal{C}^{\infty}(SM,\mathbb{R})$ be a smooth function on the unit tangent bundle $\pi:SM\to M$. We concern ourselves with the dynamical system governed by the equation
@@
This equation defines a flow $\varphi_{t}:=(\gamma(t),\dot{\gamma}(t))$ on $SM$ which reduces to the geodesic flow when $\lambda=0$. The flow models the motion of a particle under the influence of a force orthogonal to the velocity and with magnitude $\lambda$. Its generating vector field is $F=X+\lambda V$, where $X$ is the geodesic vector field on $SM$, and $V$ is the vertical vector field. The system $(M,g,\lambda)$ is called a (generalized) thermostat.
@@
If $\lambda$ does not depend on the velocity, i.e., if it corresponds to a function on $M$, then $\varphi_{t}$ is the magnetic flow associated with the magnetic field $\lambda\mu_{a}$, where $\mu_{a}$ is the area form of $(M,g)$. When $\lambda$ depends linearly on the velocity, i.e., when it corresponds to a $1$-form on $M$, we instead obtain a Gaussian thermostat, which is reversible in the sense that the flip $(x,v)\mapsto(x,-v)$ on $SM$ conjugates $\varphi_{t}$ with $\varphi_{-t}$ (just as in the case of geodesic flows). The resulting flows are interesting from a dynamical point of view because, contrary to geodesic or magnetic flows, they may not preserve any smooth volume form (see [DP07]). Gaussian thermostats also appear in geometry as the geodesic flows of certain metric connections with torsion (see [PW08]). We thus think of them as dissipative geodesic flows.
@@
We are interested in rigidity results for generalized thermostats satisfying the Anosov property. By [Ghy84, Theorem A], these flows are topologically orbit equivalent to the geodesic flow of any metric of constant negative curvature on $M$ via a Hölder homeomorphism which is in fact isotopic to the identity. In particular, this tells us that the flows of generalized thermostats are transitive and topologically mixing, so the idea is that the richness of the chaotic orbits should allow one to recover information about the system.
@@
The set up is as follows. Given two generalized thermostats $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$ on the same surface $M$, we assume there is a smooth orbit equivalence $\phi:\tilde{S}M\to SM$ which is isotopic to the identity. Here $\tilde{S}M$ is the unit tangent bundle with respect to the metric $\tilde{g}$ on $M$, and orbit equivalence means that oriented orbits are mapped to oriented orbits, i.e., there exists $c\in\mathcal{C}^{\infty}(SM,\mathbb{R}_{>0})$ such that $\phi_{*}\tilde{F}=cF$. In particular, $\phi$ is a conjugacy if $c$ is identically $1$. There is a natural identification of $SM$ with $\tilde{S}M$ by scaling the fibers via the map
$s:SM\to\tilde{S}M,\quad(x,v)\mapsto(x,v/\|v\|_{\tilde{g}}).$ (1.1)
@@
By saying that $\phi$ is isotopic to the identity we mean that $s\circ\phi:\tilde{S}M\to\tilde{S}M$ is isotopic to the identity in the usual sense.
@@
Question: If both thermostat flows are Anosov, what is the relationship, if any, between $(g,\lambda)$ and $(\tilde{g},\tilde{\lambda})$?
@@
The work in [GLP24] gives an answer in the case where $\lambda=\tilde{\lambda}=0$ and $\phi$ is a conjugacy. The metrics $g$ and $\tilde{g}$ must be isometric via an isometry isotopic to the identity. Instead of starting with a smooth conjugacy isotopic to the identity, they start with the equivalent assumption that both metrics have the same marked length spectrum. The two assumptions are also equivalent for magnetic flows, but having the same marked length spectra only guarantees a Hölder continuous conjugacy in general.
@@
More recently, progress has been made with [Reb23] to understand a deformative version of our question in the purely magnetic case, framed through the lens of marked length spectrum rigidity.
@@ 1.1. Main results
Beyond its physical motivation, the magnetic case represents the first step towards the broader goal of understanding generalized thermostats: it corresponds to the case where $\lambda=\lambda_{0}$ has Fourier degree 00 (see §2.1.4). Our first main result is the following:
@@
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$ be two Anosov magnetic systems on a smooth closed orientable surface $M$. If there is a smooth orbit equivalence isotopic to the identity between them, then there exists a smooth diffeomorphism $\psi:M\to M$, isotopic to the identity, such that $\psi^{*}\tilde{g}=e^{2f}g$ for some $f\in\mathcal{C}^{\infty}(M,\mathbb{R})$. Moreover, if the orbit equivalence is a conjugacy and $f=0$, then $\lambda=0$ if and only if $\tilde{\lambda}=0$.
We note that finding a relationship between $\lambda$ and $\tilde{\lambda}$ in the general magnetic case remains an open question. A key similarity between geodesic and magnetic flows is that they preserve the Liouville measure on $SM$. As we will explain, this allows most of the key arguments from the paper [GLP24] to also go through in the magnetic case.
For this reason, the main emphasis of this paper is instead on Gaussian thermostats. These correspond to the case where $\lambda=\lambda_{-1}+\lambda_{1}$ or, equivalently,
$\lambda=\pi_{1}^{*}\theta$ (1.2)
for some 1-form $\theta$ on $M$, where $(\pi_{1}^{*}\theta)(x,v):=\theta_{x}(v)$ denotes the restriction to $SM$ of smooth differential forms (so that we may see them as functions on $SM$). We will denote a Gaussian thermostat $(M,g,\lambda)$ by $(M,g,\theta)$ to highlight its particular form.
One can also study Gaussian thermostats using an external vector field $E$. This is the vector field on $M$ characterized by $\theta_{x}(v)=g(E(x),Jv)$, that is, the vector field dual to $\star\theta$, where $\star$ is the Hodge star operator of the metric $g$ (given by oriented rotation by $\pi/2$ in the case of $1$-forms).
As we allow $\lambda$ to have Fourier degree 1, we introduce the possibility of new dynamical features absent from the geodesic and magnetic cases. For instance, by [DP07, Theorem A], a Gaussian thermostat preserves an absolutely continuous invariant measure on $SM$ if and only if $\star\theta$ is exact. This means that the Liouville measure may no longer be preserved, and it allows for fractal SRB measures.
@@
$\mathbb{K}:=\pi^{*}K_{g}-H(\lambda)+\lambda^{2}+F(V(\lambda))$ (1.4)
used for any generalized thermostat $(M,g,\lambda)$.
@@
Let $(M,g,\theta)$ and $(M,\tilde{g},\tilde{\theta})$ be two Gaussian thermostats with $\mathbb{K},\widetilde{\mathbb{K}}<0$ on a smooth closed orientable surface $M$. If there is a smooth orbit equivalence isotopic to the identity between them, then there exists a smooth diffeomorphism $\psi:M\to M$, isotopic to the identity, such that $\psi^{*}\tilde{g}=e^{2f}g$ for some $f\in\mathcal{C}^{\infty}(M,\mathbb{R})$. Moreover, if either $\star\theta$ or ~$\tilde{\star}\tilde{\theta}$ is closed, then $\star(\psi^{*}\tilde{\theta}-\theta)$ is exact.
@@
As shown in Lemma 4.6, the scaling map defined in (1.1) yields a smooth orbit equivalence isotopic to the identity between the Gaussian thermostats $(M,g,\theta)$ and $(M,e^{2f}g,\theta+\star df)$, with a time-change by $e^{f}$. This implies that the conformal factor in our main result is optimal and that it is necessary to leave room for an exact difference when relating the $1$-forms. However, it is unclear at this stage whether the closedness condition is really necessary to establish this last relationship.
@@
Ideally, one would like to extend this result to the general Anosov case. The only place where we use the negative thermostat curvature is in showing that the Gaussian thermostats satisfy the attenuated tensor tomography problem of order $1$ (see §2.3.2). We do not have this issue in the purely magnetic case, which is why we were able to simply assume the more general Anosov property in Theorem 1.1. Removing the negative thermostat curvature assumption should also allow one to mix the magnetic case with Gaussian thermostats, i.e., to take $\lambda=\lambda_{-1}+\lambda_{0}+\lambda_{1}$.
@@
As pointed out above, there are still open questions regarding the rigidity of $\lambda$ for $\lambda$ of Fourier degree $1$. It is also unclear at this stage how much information is gained from having a genuine conjugacy versus an orbit equivalence, and whether the conjugating diffeomorphism $\phi$ itself must have some particular form as in the purely geodesic case (see [GLP24, Corollary 1.2]).
@@
After this work, a natural question is whether anything can be said for $\lambda$ of Fourier degree $\geq 2$. As we show with the no-go Lemma 2.16, the current argument does not work for these thermostats. However, there are interesting examples of such systems. For instance, when $\lambda$ is the real part of a holomorphic differential of degree $\geq 2$, the corresponding thermostat admits an interpretation as coupled vortex equations (see [MP19]). It was also shown in [MP20] that the geodesic flow of an affine connection on $M$ is, up to a time-change, the flow of a generalized thermostat with $\lambda$ of the form $\lambda=\lambda_{-3}+\lambda_{-1}+\lambda_{1}+\lambda_{3}$. Just as we have shown that a non-trivial Anosov magnetic system $(M,g,\lambda)$ cannot be smoothly conjugate to an Anosov geodesic flow $(M,g,0)$ by a conjugacy isotopic to the identity, it would be interesting to further categorize generalized thermostats.
@@
Finally, we note that Theorem A.5, which applies to generalized thermostats, was placed in the appendix to improve the overall exposition of the paper, but it represents a new result related to the injectivity of the generalized thermostat X-ray transform.
@@ 1.2. Strategy
Our main inspiration is the approach in [GLP24]. Indeed, we show that a smooth orbit equivalence isotopic to the identity determines the complex structure of the metric $g$ up to biholomorphisms isotopic to the identity (Proposition 4.2). This allows us to conclude that the two metrics $g$ and $\tilde{g}$ must be conformally equivalent via a smooth diffeomorphism of $M$ isotopic to the identity.
@@
The period matrix is defined in terms of holomorphic $1$-forms on $M$. We show with Theorem 2.15 that these can always be associated to the first Fourier modes of certain distributions $\mathcal{D}_{\text{tr},+}(SM)$ on $SM$ satisfying a transport equation and with non-negative Fourier modes (see §2.2.2). Asking for these distributions to only have non-negative Fourier modes is a critical requirement for the rest of the argument, but it does not carry over to the case of generalized thermostats when $\lambda$ has Fourier degree $\geq 2$.
We then establish in Lemma 3.6 a pairing formula showing that the integral of any holomorphic $1$-form over a thermostat geodesic $\gamma$ on $M$ (i.e., the periods of the period matrix) is the same as the integral over $\pi^{-1}(\gamma)\subseteq SM$ of an associated $2$-current invariant by $F$ and living in a certain subspace $\mathcal{F}(SM)$ (see §2.1.3). This pairing formula then tells us that the smooth orbit equivalence preserves the period matrix.
At a high level, there are two main challenges and departures from [GLP24]: the first is in handling a general orbit equivalence instead of a conjugacy, and the second is in dealing with the fact that Gaussian thermostats may not be volume-preserving.
The presence of a non-zero divergence with respect to the Liouville form manifests itself in a few ways. First, instead of flow-invariant distributions, the right object of study becomes solutions to the dual transport equation. This subspace is no longer preserved by the pullback of the orbit equivalence, so we have to introduce the space $\mathcal{F}(SM)$ of $2$-currents mentioned above and establish a one-to-one relationship with the distributions solving the transport equation (Lemma 2.3). We then have to check that the wavefront set analysis is unaffected by factoring the correspondence through this space (Lemma 2.4) and that $\mathcal{D}_{\text{tr},+}(SM)$ is mapped to $\mathcal{D}_{\text{tr},+}(\tilde{S}M)$ (Proposition 3.3).
@@
For the pairing formula previously described, we have replaced the role of the Liouville form with that of a certain form defined in (2.2). Finally, to relate $\lambda$ with $\tilde{\lambda}$, we rely on new arguments which at their core involve the smooth Livšic theorem.

### v2

@@ Abstract.
Let $M$ be a smooth closed oriented surface. Gaussian thermostats on $M$ correspond to the geodesic flows arising from metric connections, including those with non-zero torsion. These flows may not preserve any absolutely continuous measure. We prove that if two Gaussian thermostats on $M$ with negative thermostat curvature are related by a smooth orbit equivalence isotopic to the identity, then the two background metrics are conformally equivalent via a smooth diffeomorphism of $M$ isotopic to the identity. We also give a relationship between the thermostat forms themselves. Finally, we prove the same result for Anosov magnetic flows.
@@ 1. Introduction
Let $(M,g)$ be a closed oriented Riemannian surface and let $\lambda\in\mathcal{C}^{\infty}(SM,\mathbb{R})$ be a smooth function on the unit tangent bundle $\pi:SM\to M$. We concern ourselves with the dynamical system governed by the equation
@@
This equation defines a flow $\varphi_{t}(\gamma(0),\dot{\gamma}(0))\coloneqq(\gamma(t),\dot{\gamma}(t))$ on $SM$ which reduces to the geodesic flow when $\lambda=0$. The flow models the motion of a particle under the influence of a force orthogonal to the velocity and with magnitude $\lambda$. Its generating vector field is $F=X+\lambda V$, where $X$ is the geodesic vector field on $SM$ and $V$ is the vertical vector field. The system $(M,g,\lambda)$ is called a thermostat.
@@
If $\lambda$ does not depend on velocity, that is, if it corresponds to a function on $M$, then $\varphi_{t}$ is the magnetic flow associated with the magnetic field $\lambda\mu_{a}$, where $\mu_{a}$ is the area form of $(M,g)$. When $\lambda$ depends linearly on velocity, that is, when it corresponds to a $1$-form on $M$, we instead obtain a Gaussian thermostat, which is reversible in the sense that the flip $(x,v)\mapsto(x,-v)$ on $SM$ conjugates $\varphi_{t}$ with $\varphi_{-t}$ (just as in the case of geodesic flows). The resulting flows are interesting from a dynamical point of view because, in contrast to geodesic or magnetic flows, they may not preserve any absolutely continuous measure (see [DP07]). Gaussian thermostats also appear in geometry as the geodesic flows of metric connections, including those with non-zero torsion (see [PW08]). We thus think of them as dissipative geodesic flows.
@@
We are interested in rigidity results for thermostats satisfying the Anosov property. By [Ghy84, Theorem A], these flows are topologically orbit equivalent to the geodesic flow of any metric with constant negative curvature on $M$ via a Hölder homeomorphism which is in fact isotopic to the identity. In particular, this tells us that the flows of thermostats are transitive and topologically mixing, so the idea is that the richness of the chaotic orbits should allow one to recover information about the system.
@@
The set-up is as follows. Given two thermostats $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$ on the same surface $M$, we assume there is a smooth orbit equivalence $\phi:\widetilde{S}M\to SM$ which is isotopic to the identity. Here, $\widetilde{S}M$ is the unit tangent bundle with respect to the metric $\tilde{g}$ on $M$, and orbit equivalence means that oriented orbits are mapped to oriented orbits, that is, there exists $c\in\mathcal{C}^{\infty}(SM,\mathbb{R}_{>0})$ such that $\phi_{*}\widetilde{F}=cF$. In particular, $\phi$ is a conjugacy if $c$ is identically $1$. There is a natural identification of $SM$ with $\widetilde{S}M$ by scaling the fibres via the map
$s:SM\to\widetilde{S}M,\qquad s(x,v)\coloneqq(x,v/\|v\|_{\tilde{g}}).$ (1.1)
@@
By saying that $\phi$ is isotopic to the identity, we mean that $s\circ\phi:\widetilde{S}M\to\widetilde{S}M$ is isotopic to the identity in the usual sense.
Question 1.1.
@@
If both thermostat flows are Anosov, what is the relationship, if any, between $(g,\lambda)$ and $(\tilde{g},\tilde{\lambda})$?
@@
The work in [GLP25] gives an answer when $\lambda=\tilde{\lambda}=0$ and $\phi$ is a conjugacy: $g$ and $\tilde{g}$ must be isometric via an isometry isotopic to the identity. Rather than assuming a smooth conjugacy isotopic to the identity, they work with the equivalent condition that the metrics have the same marked length spectrum. This equivalence also holds for magnetic flows (using [GR23, Theorem 1.1]). However, while the notion of marked length spectrum still makes sense for any thermostat (each non-trivial free homotopy class on $M$ contains a unique closed thermostat geodesic by [Ghy84, Theorem A]), equality of marked length spectra only guarantees a Hölder continuous conjugacy.
@@
More recently, progress has been made in [Mar23, Mar24] to understand a deformative version of our question in the purely magnetic case, framed through the lens of marked length spectrum rigidity.
@@ 1.1. Main results
Beyond its physical motivation, the magnetic case represents the first step towards the broader goal of understanding thermostats: it corresponds to the case where $\lambda=\lambda_{0}$ has Fourier degree 00 (see §2.1.4). Our first result is the following theorem.
@@
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$ be two Anosov magnetic systems. If there is a smooth orbit equivalence isotopic to the identity between them, then there exists a smooth diffeomorphism $\psi:M\to M$ isotopic to the identity such that $\psi^{*}\tilde{g}=e^{2f}g$ for some $f\in\mathcal{C}^{\infty}(M,\mathbb{R})$. Moreover, if the orbit equivalence is a conjugacy and $f=0$, then $\lambda=0$ if and only if $\tilde{\lambda}=0$.
We note that finding a relationship between $\lambda$ and $\tilde{\lambda}$ in the general magnetic case remains an open question. A key similarity between geodesic and magnetic flows is that they preserve the Liouville measure on $SM$. As we will explain, this allows most of the key arguments from the paper [GLP25] to also go through in the magnetic case.
For this reason, the main emphasis of this paper is instead on Gaussian thermostats. These correspond to the case where $\lambda=\lambda_{-1}+\lambda_{1}$ or, equivalently, $\lambda=\ell_{1}\theta$ for some 1-form $\theta$ on $M$, where
$(\ell_{1}\theta)(x,v)\coloneqq\theta_{x}(v)$ (1.2)
denotes the restriction to $SM$ of smooth differential forms (so that we may see them as functions on $SM$). We will denote a Gaussian thermostat $(M,g,\lambda)$ by $(M,g,\theta)$ to highlight its particular form.
One can also study Gaussian thermostats using an external vector field $E$. This is the vector field on $M$ uniquely characterized by $\theta_{x}(v)=g_{x}(E(x),Jv)$, that is, the vector field dual to $\star\theta$, where $\star$ is the Hodge star operator of the metric $g$.
As we allow $\lambda$ to have Fourier degree 1, we introduce the possibility of new dynamical features absent from the geodesic and magnetic cases. For instance, by [DP07, Theorem A], a Gaussian thermostat preserves an absolutely continuous invariant measure on $SM$ if and only if $\star\theta$ is exact. This means in particular that the Liouville measure may no longer be preserved and it allows for fractal Sinai–Ruelle–Bowen measures.
@@
$\mathbb{K}\coloneqq\pi^{*}K_{g}-H\lambda+\lambda^{2}+FV\lambda$ (1.4)
used for any thermostat $(M,g,\lambda)$.
@@
Let $(M,g,\theta)$ and $(M,\tilde{g},\tilde{\theta})$ be two Gaussian thermostats with $\mathbb{K},\widetilde{\mathbb{K}}<0$. If there is a smooth orbit equivalence isotopic to the identity between them, then there is a smooth diffeomorphism $\psi:M\to M$ isotopic to the identity such that $\psi^{*}\tilde{g}=e^{2f}g$ for some $f\in\mathcal{C}^{\infty}(M,\mathbb{R})$. Moreover, if $\star\theta$ or ~$\tilde{\star}\tilde{\theta}$ is closed, then $\star(\psi^{*}\tilde{\theta}-\theta)$ is exact.
@@
As shown in Lemma 4.6, the scaling map defined in equation (1.1) yields a smooth orbit equivalence isotopic to the identity between the Gaussian thermostats $(M,g,\theta)$ and $(M,e^{2f}g,\theta+\star df)$, with a time-change by $e^{f}$. This implies that the conformal factor in our main result is optimal and that it is necessary to leave room for an exact difference when relating the $1$-forms. However, it is unclear at this stage whether the closedness condition is really necessary to establish this last relationship.
@@
Ideally, one would like to extend this result to the general Anosov case. The only place where we use the negative thermostat curvature is in showing that the Gaussian thermostats satisfy the attenuated tensor tomography problem of order $1$ (see §2.3.2). We do not have this issue in the purely magnetic case, which is why we were able to simply assume the more general Anosov property in Theorem 1.2. Removing the negative thermostat curvature assumption should also allow one to mix the magnetic case with Gaussian thermostats, that is, to take $\lambda=\lambda_{-1}+\lambda_{0}+\lambda_{1}$.
@@
As pointed out above, there are still open questions regarding the rigidity of $\lambda$ for $\lambda$ of Fourier degree $1$. It is also unclear at this stage how much information is gained from having a genuine conjugacy versus an orbit equivalence, and whether the conjugating diffeomorphism $\phi$ itself must have some particular form as in the purely geodesic case (see [GLP25, Corollary 1.3]).
@@
After this work, a natural question is whether anything can be said for $\lambda$ of Fourier degree $\geq 2$. As we show with the no-go Lemma 2.16, the current argument does not work for these thermostats. However, there are interesting examples of such systems. For instance, when $\lambda$ is the real part of a holomorphic differential of degree $\geq 2$, the corresponding thermostat admits an interpretation as coupled vortex equations (see [MP19, MP22]). It was also shown in [MP20] that the geodesic flow of an affine connection on $M$ is, up to a time-change, the flow of a thermostat with $\lambda$ of the form $\lambda=\lambda_{-3}+\lambda_{-1}+\lambda_{1}+\lambda_{3}$. Just as Theorem 1.2 implies that an Anosov magnetic system $(M,g,\lambda)$ with $\lambda\neq 0$ cannot be smoothly conjugate to an Anosov geodesic flow $(M,g,0)$ with the same metric by a conjugacy isotopic to the identity, it would be interesting to further categorize thermostats.
@@
Finally, we note that Theorem A.5, which applies to thermostats, was placed in the appendix to improve the overall exposition of the paper, but it represents a new result related to the injectivity of the thermostat X-ray transform.
@@ 1.2. Strategy
Our main inspiration is the approach in [GLP25]. Indeed, we show that a smooth orbit equivalence isotopic to the identity determines the complex structure of the metric $g$ up to biholomorphisms isotopic to the identity (Proposition 4.2). This allows us to conclude that the two metrics $g$ and $\tilde{g}$ must be conformally equivalent via a smooth diffeomorphism of $M$ isotopic to the identity.
@@
The period matrix is defined in terms of holomorphic $1$-forms on $M$. We show with Theorem 2.15 that these can always be associated to the first Fourier modes of certain distributions $\mathcal{D}_{\text{tr},+}(SM)$ on $SM$ satisfying a transport equation and with non-negative Fourier modes (see §2.2.2). Asking for these distributions to only have non-negative Fourier modes is a critical requirement for the rest of the argument, but it does not carry over to the case of thermostats when $\lambda$ has Fourier degree $\geq 2$.
We then establish in Lemma 3.6 a pairing formula showing that the integral of any holomorphic $1$-form over a thermostat geodesic $\gamma$ on $M$ (that is, the periods of the period matrix) is the same as the integral over $\pi^{-1}(\gamma)\subset SM$ of an associated $2$-current invariant by $F$ and living in a certain subspace $\mathcal{F}(SM)$ (see §2.1.3). This pairing formula then tells us that the smooth orbit equivalence preserves the period matrix.
At a high level, there are two main challenges and departures from [GLP25]: the first is in handling a general orbit equivalence instead of a conjugacy, and the second is in dealing with the fact that Gaussian thermostats may not be volume-preserving.
The presence of a non-zero divergence with respect to the Liouville volume form manifests in several ways. First, instead of flow-invariant distributions, the right object of study becomes solutions to the dual transport equation. This subspace is no longer preserved by the pullback of the orbit equivalence, so we have to introduce the space $\mathcal{F}(SM)$ of $2$-currents mentioned above and establish a one-to-one relationship with the distributions solving the transport equation (Lemma 2.3). We then have to check that the wavefront set analysis is unaffected by factoring the correspondence through this space (Lemma 2.4) and that $\mathcal{D}_{\text{tr},+}(SM)$ is mapped to $\mathcal{D}_{\text{tr},+}(\widetilde{S}M)$ (Proposition 3.3).
@@
For the pairing formula previously described, we have replaced the role of the Liouville form with that of a certain form defined in equation (2.2). Finally, to relate $\lambda$ with $\tilde{\lambda}$, we rely on new arguments which, at their core, involve the smooth Livšic theorem.